OpenStudy (anonymous):

Write a rule for the linear function in the table. x f(x) 1 –7 2 –10 3 –13 4 –16 A. f(x) = -1/3x – 4 B. f(x) = 3x + 4 C. f(x) = –3x – 4 D. f(x) = x – 7

OpenStudy (zale101):

TO write a function rule from a table is by looking for a pattern. The formula is y=mx+b in function notation f(x)=mx+b The y-intercept (b) is thevalue of f(x) when x=0

OpenStudy (anonymous):

oh ok wat next

OpenStudy (anonymous):

anyone help me please

OpenStudy (zale101):

(1,-7); (2,-10) y+7= (-10+7)/1 (x-1) y=-x+1-7 f(x)=-x-6

OpenStudy (anonymous):

its c

OpenStudy (zale101):

C f(x)=-3x-4 is the right answer

OpenStudy (anonymous):

yeah